OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (anonymous):

if \(\large u =\langle u_1, u_2 \rangle\) then \(\large \frac{1}{|u|}\langle u_1, u_2\rangle\) is the unit vector in the direction of u :) Just multiply the vector with the scalar 1/|u|

OpenStudy (anonymous):

<-4/5, -3/5>?

OpenStudy (anonymous):

Looks correct to me :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

You can always check if it's an unit vector by calculating the magnitude: sqrt((-4/5)^2 + (-3/5)^2) = sqrt(16/25 + 9/25) = sqrt(1) = 1